# On quantum learning algorithms for noisy linear problems

### v1

@@ 5.3 Lattice reduction
Given two (measured) resized samples $(a_{j,1}^{\prime},b_{j,1}^{\prime}=a_{j,1}^{\prime}s_{j}+e_{j,1}^{\prime}\bmod q)$, $(a_{j,2}^{\prime},b_{j,2}^{\prime}=a_{j,2}^{\prime}s_{j}+e_{j,2}^{\prime}\bmod q)$, let us consider the following lattice:
$\mathcal{L}$ $=$ $\big{\{}{\bm{v}}\in\mathbb{Z}^{2}:{\bm{v}}\equiv x{\bm{a}}\bmod q\textup{ for some }x\in\mathbb{Z}\big{\}}$
@@
where ${\bm{a}}=(a_{j,1}^{\prime},a_{j,2}^{\prime})$ and ‘$\bmod~{}q$’ applies to each component. Assume that $a_{j,1}^{\prime}$ is invertible in $\mathbb{Z}_{q}$, then $(1,(a_{j,1}^{\prime})^{-1}a_{j,2}\bmod q),(0,q)$ is a basis of $\mathcal{L}$ and so $\dim(\mathcal{L})=2$, $\det(\mathcal{L})=q$.
The purpose of introducing $\mathcal{L}$ is to reduce finding $s_{j}$ to the closest vector problem (CVP). Notice that if ${\bm{e}}:=(e_{j,1}^{\prime},e_{j,2}^{\prime})={\bm{0}}$, then ${\bm{b}}:=(b_{j,1}^{\prime},b_{j,2}^{\prime})\in\mathcal{L}$. For nonzero small $\|{\bm{e}}\|$, we thus expect that ${\bm{b}}\equiv s_{j}{\bm{a}}+{\bm{e}}\bmod q$ is close to a vector $s_{j}{\bm{a}}\bmod q\in\mathcal{L}$. Therefore the task is to solve CVP in $\mathcal{L}$ for target ${\bm{b}}$.
Generally the CVP is taken to be a hard problem, but in two dimension there exists an efficient solution: Lagrange-Gauss reduction is first applied to find a good basis of $\mathcal{L}$, and then Babai’s nearest plane algorithm is used to find the closest vector in $\mathcal{L}$ to ${\bm{b}}$ [3].
By applying the Lagrange-Gauss reduction algorithm, one can compute basis vectors ${\bm{g}}_{1},{\bm{g}}_{2}\in\mathcal{L}$ such that $\|{\bm{g}}_{i}\|=\lambda_{i}(\mathcal{L})$ in time $O(\log^{3}q)$ [11, Chapter 17], where $\lambda_{i}(\mathcal{L})$ is the $i$-th successive minimum of $\mathcal{L}$. For given basis vectors $\mathbf{b}_{1},\ldots,\mathbf{b}_{\dim(\mathcal{L})}$ and a target ${\bm{u}}$, if there is a vector ${\bm{v}}\in\mathcal{L}$ satisfying $\|{\bm{v}}-{\bm{u}}\|\leq\frac{1}{2}\min_{i\in\{1,\ldots,\dim(\mathcal{L})\}}\{\|\mathbf{b}_{i}^{*}\|\}$, then the Babai’s nearest plane algorithm always outputs the vector ${\bm{v}}$ in time $O(\log^{2}q)$ [11, Chapter 18], where $\mathbf{b}_{i}^{*}$ is the $i$-th Gram-Schmidt orthogonalized vector. Assuming the Gaussian heuristic [11, Chapter 16] holds, $\|{\bm{g}}_{1}\|\approx\sqrt{\det(\mathcal{L})}=\sqrt{q}$. Since ${\bm{g}}_{1}={\bm{g}}_{1}^{*}$ and $\|{\bm{g}}_{1}^{*}\|\|{\bm{g}}_{2}^{*}\|=\det(\mathcal{L})$, we have $\|{\bm{g}}_{2}^{*}\|\approx\sqrt{q}$. Therefore Babai’s algorithm results in a vector ${\bm{v}}\in\mathcal{L}$ such that $\|{\bm{v}}-{\bm{b}}\|\leq\frac{1}{2}\sqrt{q}$. If $\|{\bm{e}}\|\leq\frac{1}{2}\sqrt{q}$, then one would expect ${\bm{v}}\equiv s_{j}{\bm{a}}\bmod q$. Indeed, since $|e_{j,i}^{\prime}|\leq k^{\prime}$ (i.e., $\|{\bm{e}}\|\leq\sqrt{2}k^{\prime}$), a condition for the above strategy to work is
$8k^{\prime 2}\leq q\enspace.$ (11)
Equation (6) in [19] assumes $k^{\prime}\ll q$, thus Eq. (11) likely hold.

### v2

@@ Introduction
We are now ready to describe the method. Suppose we are given two classical samples $(a_{j,k}^{\prime},b_{j,k}^{\prime}=a_{j,k}^{\prime}s_{j}+e_{j,k}^{\prime}\bmod q)$ for $k=1,2$, by measuring the resized quantum samples. Assume that $a_{j,1}^{\prime}$ is invertible in $\mathbb{Z}_{q}$. Let ${\bm{a}}=(a_{j,1}^{\prime},a_{j,2}^{\prime})$ and consider the following lattice:
=$\mathcal{L}=\big{\{}{\bm{v}}\in\mathbb{Z}^{2}:{\bm{v}}\equiv x{\bm{a}}\bmod q\textup{ for some }x\in\mathbb{Z}\big{\}}\enspace,$
@@
where $\bmod\leavevmode\nobreak\ q$ applies to each component. Notice that by letting ${\bm{v}}=(v_{1},v_{2})\equiv x{\bm{a}}$ for some $x\in\mathbb{Z}$, we have $x\equiv(a_{j,1}^{\prime})^{-1}v_{1}\bmod q$. This shows that two vectors $\mathbf{b}_{1}=(1,(a_{j,1}^{\prime})^{-1}a_{j,2}^{\prime}\bmod q)$ and $\mathbf{b}_{2}=(0,q)$ form a basis of $\mathcal{L}$, that is,
=$\mathcal{L}=\{{\bm{v}}:{\bm{v}}=z_{1}\mathbf{b}_{1}+z_{2}\mathbf{b}_{2},\leavevmode\nobreak\ z_{1},z_{2}\in\mathbb{Z}\},$
and thus $\dim(\mathcal{L})=2$, $\det(\mathcal{L})=\det\!\big{(}\big{[}{\mathbf{b}_{1}\atop\mathbf{b}_{2}}\big{]}\big{)}=q$. As explained earlier, the purpose of introducing $\mathcal{L}$ is to reduce finding $s_{j}$ to solving CVP: if one manages to find ${\bm{v}}\in\mathcal{L}$ that is close to ${\bm{b}}=(b_{j,1}^{\prime},b_{j,2}^{\prime})$, then the error is obtained by ${\bm{e}}={\bm{b}}-{\bm{v}}$.
By using Lagrange-Gauss algorithm with inputs $\mathbf{b}_{1},\mathbf{b}_{2}$, we have basis vectors ${\bm{g}}_{1},{\bm{g}}_{2}\in\mathcal{L}$ such that $\|{\bm{g}}_{i}\|=\lambda_{i}(\mathcal{L})$ for $i=1,2$. Assuming the Gaussian heuristic holds, $\lambda_{1}(\mathcal{L})=\|{\bm{g}}_{1}\|\approx\sqrt{\det(\mathcal{L})}=\sqrt{q}$. Let $\mathbf{g}_{1}^{*},\mathbf{g}_{2}^{*}$ be Gram-Schmidt orthogonalized vectors of ${\bm{g}}_{1},{\bm{g}}_{2}$. Due to the properties of Gram-Schmidt orthogonalized vectors, we know that ${\bm{g}}_{1}=\mathbf{g}_{1}^{*}$ and $\det(\mathcal{L})=\det\!\big{(}\big{[}{\mathbf{b}_{1}\atop\mathbf{b}_{2}}\big{]}\big{)}=\det\!\big{(}\big{[}{\mathbf{g}_{1}^{\ast}\atop\mathbf{g}_{2}^{\ast}}\big{]}\big{)}=\|\mathbf{g}_{1}^{*}\|\|\mathbf{g}_{2}^{*}\|$. Therefore, $\|\mathbf{g}_{2}^{*}\|\approx\sqrt{q}$.
Since ${\bm{b}}\equiv s_{j}{\bm{a}}+{\bm{e}}\bmod q$, we have ${\bm{b}}=s_{j}{\bm{a}}+{\bm{e}}+q{\bm{l}}$ for some ${\bm{l}}\in\mathbb{Z}^{2}$ and $\|{\bm{b}}-(s_{j}{\bm{a}}+q{\bm{l}})\|=\|{\bm{e}}\|$. Thus, if $\|{\bm{e}}\|<\frac{1}{2}\min\{\|\mathbf{g}_{1}^{\ast}\|,\|\mathbf{g}_{2}^{\ast}\|\}\approx\frac{1}{2}\sqrt{q}$, then taking ${\bm{g}}_{1},{\bm{g}}_{2}$ and ${\bm{b}}$ as inputs, Babai’s algorithm outputs the vector ${\bm{w}}=s_{j}{\bm{a}}+q{\bm{l}}$ by Proposition 1. Computing ${\bm{w}}\bmod q$ then yields $s_{j}{\bm{a}}\bmod q$, from which we recover $s_{j}$. Indeed, since $|e_{j,i}^{\prime}|\leq\xi^{\prime}$ (i.e., $\|{\bm{e}}\|\leq\sqrt{2}\xi^{\prime}$), a condition for the algorithm to work is
$8\xi^{\prime 2}\lesssim q\enspace.$ (8)
In the previous paper, the authors assume $\xi^{\prime}\ll q$, thus Equation (8) likely holds. Considering the complexity of each subroutine algorithm, the total running time is $\mathsf{poly}(\log(q))$.
Figure 2: Success rate of the proposed algorithm as a function of the error bound for (a) $q=3329$ (Kyber) and for (b) $q=2^{23}-2^{13}+1$ (Dilithium). In each figure, the condition given by Equation (8) is around 20 and 1023, respectively. In (b), the interval in the horizontal axis is 10.
The above analysis relies on Gaussian heuristic which is not guaranteed to work beneficially on every instance. To verify the validity of the approach, numerical tests are carried out summarized in Figure 2. Each subfigure shows the success rate of the lattice reduction method for chosen modulus $q$ from Kyber [22] and Dilithium [23], respectively. For each error bound $\xi^{\prime}$, the following procedure is conducted 1000 times: a pair of resized samples is produced randomly (including the secret) and the algorithm is applied to output a candidate for the secret. The success rate is the number of times the algorithm correctly outputs the secret divided by 1000. As can be seen from the figure, the algorithm succeeds better as the bound gets smaller. There is a chance that the algorithm fails for small $\|{\bm{e}}\|$ due to unusually short Gram-Schmidt vectors, but such cases can be overcome by getting another sample and trying a different lattice.
Conclusion
A definitive conclusion on whether the linear learning algorithms will be immediately useful once a powerful quantum computer is ready is not yet drawn from this work, but we have shown that the RLWE problem can be tackled if LWE is indeed solvable quantumly. In addition, it is also shown that assumptions taken in solving SIS and size-reduced LWE problems lead to efficient classical algorithms, implying that the previously claimed quantum advantages need to be reexamined. A further multidisciplinary study on the subject may estimate the true feasibility of the algorithms.
Data availability
All data generated or analysed during this study are included in this published article.